OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4.

TheSmartOne (thesmartone):

Hi, welcome to OpenStudy! @latania22 Distribute the -2 on both sides of your question. What do you get?

TheSmartOne (thesmartone):

equation* subconsciously thinking equation but I write question lol :p

OpenStudy (anonymous):

-2(x+3)=-2(x+1)-4 =-2x-6=-2x-2-4 (Opening the brackets) @TheSmartOne :^ The equation seems inconsistent, doesn't it?

TheSmartOne (thesmartone):

No, it has infinitely many solutions so it can't be inconsistent.

OpenStudy (anonymous):

So, family of lines?